rewrite the equation in function form:
-x +y = 6

Answer by rapaljer(4551)

Solve for y:
-x + y = 6

Add +x to each side of the equation:
-x+x +y = x+6
y= x+ 6

y = f(x), so
y= f(x) = x+ 6

R^2 at SCC